The measure of an angle exceeds twice the measure of the complement of the angle by 20. What is the measure of half of the complement?
Answer by mananth(16946)

let the angle be x
the complement = 90-x
x=2(90-x)+20
x=180-2x+20
2x+x=200
3x=200
x=66.66
complement = 23.34
half of complement = 11.67 deg
..
m.ananth@hotmail.ca